# A Style-Based Profiling Framework for Quantifying the Synthetic-to-Real Gap in Autonomous Driving Datasets

### v2

@@ Abstract
Ensuring the reliability of autonomous driving perception systems requires extensive environment-based testing, yet real-world execution is often impractical. Synthetic datasets have therefore emerged as a promising alternative, offering advantages such as cost-effectiveness, bias free labeling, and controllable scenarios. However, the domain gap between synthetic and real-world datasets remains a major obstacle to model generalization. To address this challenge from a data-centric perspective, this paper introduces a profile extraction and discovery framework for characterizing the style profiles underlying both synthetic and real image datasets. We propose Style Embedding Distribution Discrepancy (SEDD) as a novel evaluation metric. Our framework combines Gram matrix-based style extraction with metric learning optimized for intra-class compactness and inter-class separation to extract style embeddings. Furthermore, we establish a benchmark using publicly available datasets. Experiments are conducted on a variety of datasets and sim-to-real methods, and the results show that our method is capable of quantifying the synthetic-to-real gap. This work provides a standardized profiling-based quality control paradigm that enables systematic diagnosis and targeted enhancement of synthetic datasets, advancing future development of data-driven autonomous driving systems.
@@ I Introduction
To address these issues, we introduce a novel style-based profiling framework with a metric, Style Embedding Distribution Discrepancy (SEDD). The main contributions of this work are as follows:
  • We innovatively model the discrepancies between synthetic and real data as differences in extracted style profiles and propose a data fidelity metric Style Embedding Distribution Discrepancy (SEDD).
  • We utilize the publicly available real datasets and synthetic datasets to establish a benchmark for evaluating the synthetic-to-real gap.
  • A large number of experiments have been conducted on real datasets, synthetic datasets to demonstrate the effectiveness of our proposed profiling framework.
  • We apply the proposed profiling framework to sim-to-real approaches in autonomous driving, providing quantitative metrics and visualization tools for their evaluation.
@@ II-A NR-IQA for Single Images
Since many synthetic datasets lack real-world counterparts [9, 5], it is necessary to employ No-Reference Image Quality Assessment (NR-IQA) methods. Classical approaches such as BRISQUE [19] rely on natural scene statistics with SVM regression. With the advent of deep learning, data-driven NR-IQA methods have have emerged, including NIMA [20], which employs end-to-end CNN training, and Unique [21], which leverages samples from multiple databases to achieve robust generalization.
@@ IV-A1 Datasets
Real-world datasets KITTI [22], Cityscapes [27] and synthetic datasets Virtual KITTI [7], Virtual KITTI 2 [8] are used to train our reference benchmark. For synthetic datasets, weather conditions, ‘clone’, ‘morning’ and ‘overcast’, are selected. The data is partitioned into training, validation, and test sets with a ratio of 6:2:2. Additionally, we evaluate the sim-to-real methods. The photorealism enhance series include Colortransfer [23], CUT [24], EPE [25]. CARLA2Real [26] series consist of CARLA2KITTI and CARLA2CITY.
@@ IV-B1 Quantification Results on Validation Set
IV-B2 Compared with NR-IQA
The results are summarized in Table II, where the best performance is highlighted in bold and the second-best is underlined. The results demonstrate that NR-IQA methods do not sort correctly, while only our method correctly identifies KITTI as the most realistic and Virtual KITTI as the least realistic.
@@ IV-B2 Compared with NR-IQA
Datasets
Brisque
$\downarrow$ [19]
NIMA
$\uparrow$ [20]
Unique
$\uparrow$ [21]
$\mathit{SEDD}_{1}$ $\downarrow$ $\mathit{SEDD}_{2}$ $\downarrow$
@@ IV-B3 Quantification Results on Test Set
The results of our model on test set are shown in Table III. ‘-’ indicates that the experiment was not performed. The $\mathit{SEDD}_{1}$ and $\mathit{SEDD}_{2}$ of the real dataset are all smaller than those of the synthetic dataset. The visualization results on the real datasets are shown in Fig. 4b and results on the synthetic datasets are shown in Fig. 4c. Unlike other methods that have not been tested on both Virtual KITTI and Virtual KITTI 2, two datasets with extremely similar scenarios, our approach is capable of making accurate distinctions and aligns well with subjective visual assessments. Additionally, other methods rely solely on datasets captured under a single weather condition. In contrast, our method can make an effective distinction and is consistent with subjective visualization across diverse conditions.
IV-B4 Sim-to-real Gap Quantification
The results for sim-to-real methods are presented in Table IV. The smaller the value, the more effective the sim-to-real method. Visualization results for photorealism enhance series are shown in Fig. 4d and for CARLA2Real series are shown in Fig. 4e and Fig. 4f. The enhanced images exhibit lower $\mathit{SEDD}_{1}$ and $\mathit{SEDD}_{2}$ values, indicating that our proposed method effectively quantifies the impact of sim-to-real translation.
@@ IV-B4 Sim-to-real Gap Quantification
Photo-
realism
Enhance
Colortransfer[23] 0.242 0.384
@@
CARLA
2Real
Raw(CARLA)[47] 0.193 0.307
@@ V Conclusion
In this work, we propose a novel style-based profiling framework with a metric, Style Embedding Distribution Discrepancy (SEDD). Using a ResNet-based feature extractor, our approach derives style embeddings through Gram matrices and optimizes them with metric learning, enabling similarity quantification between synthetic and real datasets. Experiments validate the effectiveness of the proposed metric. This work transforms the synthetic-to-real gap into quantifiable objective indicators, eliminating reliance on subjective experience. It provides a standardized profiling-based tool for the quality control of synthetic datasets in autonomous driving, facilitating more targeted improvements. Future work could extend this framework to generative model-based synthetic datasets, such as world model related work.

### v3

@@ Abstract
Ensuring the reliability of autonomous driving perception systems requires extensive environment-based testing, yet real-world execution is often impractical. Synthetic datasets have therefore emerged as a promising alternative, offering advantages such as cost-effectiveness, bias free labeling, and controllable scenarios. However, the domain gap between synthetic and real-world datasets remains a major obstacle to model generalization. To address this challenge from a data-centric perspective, this paper introduces a profile extraction and discovery framework for characterizing the style profiles underlying both synthetic and real image datasets. We propose Style Embedding Distribution Discrepancy (SEDD) as a novel evaluation metric. Our framework combines Gram matrix-based style extraction with metric learning optimized for intra-class compactness and inter-class separation to extract style embeddings. Furthermore, we establish a benchmark using publicly available datasets. Extensive experiments demonstrate that SEDD aligns with human perception where traditional NR-IQA metrics fail. Specifically, it correctly identifies the superior fidelity of Virtual KITTI 2 over Virtual KITTI and quantifies the gap reduction achieved by sim-to-real methods. This work provides a standardized proactive quality control paradigm that enables the systematic diagnosis of dataset deficiencies and guides the selection of optimal sim-to-real adaptation strategies, advancing future development of data-driven autonomous driving systems.
©2026 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ I Introduction
To address these issues, we introduce a novel style-based profiling framework with a metric, Style Embedding Distribution Discrepancy (SEDD). SEDD functions as a proactive diagnostic tool within the data generation pipeline. It allows researchers to screen environmental parameters to minimize the distributional distance to the real domain before large-scale rendering and aids in model selection for domain adaptation by identifying which translation method yields the most realistic style. The main contributions of this work are as follows:
  • We innovatively model the discrepancies between synthetic and real data as differences in extracted style and propose a data fidelity metric SEDD.
  • We utilize the publicly available datasets to establish a benchmark for evaluating the synthetic-to-real gap.
  • Experiment results demonstrate that SEDD outperforms traditional metrics by accurately ranking dataset fidelity.
  • We apply the proposed profiling framework to sim-to-real approaches, providing quantitative metrics and visualization tools that effectively validate the gap reduction achieved by various photorealism enhancement methods.
@@ II-A NR-IQA for Single Images
Since many synthetic datasets lack real-world counterparts [36, 30], it is necessary to employ No-Reference Image Quality Assessment (NR-IQA) methods. Classical approaches such as BRISQUE [23] rely on natural scene statistics with SVM regression. With the advent of deep learning, data-driven NR-IQA methods have have emerged, including NIMA [37], which employs end-to-end CNN training, and Unique [46], which leverages samples from multiple databases to generalize.
@@ IV-A1 Datasets
Real-world datasets KITTI [13], Cityscapes [4] and synthetic datasets Virtual KITTI [10], Virtual KITTI 2 [1] are used to train our reference benchmark. For synthetic datasets, weather conditions, ‘clone’, ‘morning’ and ‘overcast’, are selected. The data is partitioned into training, validation, and test sets with a ratio of 6:2:2. Additionally, we evaluate the sim-to-real methods. The photorealism enhance series include Colortransfer [28], CUT [26], EPE [29]. CARLA2Real [27] series consist of CARLA2KITTI and CARLA2CITY. For the nuScenes dataset, only the front camera images are employed. Regarding the SHIFT dataset, images set to ‘continuous’ ‘10x’ ‘front’ are selected. Additionally, night-time samples are manually removed from both the nuScenes and SHIFT datasets. In training, the KITTI series of images included in each batch are specifically counterparts, i.e., the seven images are from the same moment in the same sequence.
@@ IV-B1 Quantification Results on Validation Set
Datasets Brisque $\downarrow$ [23] NIMA $\uparrow$ [37] Unique $\uparrow$ [46] $\mathit{SEDD}_{1}$ $\downarrow$ $\mathit{SEDD}_{2}$ $\downarrow$
@@
IV-B2 Compared with NR-IQA
The results are summarized in Table II, where the best performance is highlighted in bold and the second-best is underlined. The results demonstrate that NR-IQA methods do not sort correctly, while only our method correctly identifies KITTI as the most realistic and Virtual KITTI as the least realistic. This comparison highlights a fundamental advantage of our approach: unlike NR-IQA metrics that prioritize perceptual image quality, often erroneously assigning higher scores to clear but unrealistic synthetic images, SEDD explicitly measures the distributional deviation from the real domain style. This ensures a fidelity ranking that aligns with objective realism rather than mere visual clarity.
@@ IV-B3 Quantification Results on Test Set
The results of our model on test set are shown in Table III. ‘-’ indicates that the experiment was not performed. This absence stems from the lack of open-source implementations for certain baselines, which prevented a complete comparative evaluation. The $\mathit{SEDD}_{1}$ and $\mathit{SEDD}_{2}$ of the real dataset are all smaller than those of the synthetic dataset. The visualization results on the real datasets are shown in Fig. 4b and results on the synthetic datasets are shown in Fig. 4c. Unlike other methods that have not been tested on both Virtual KITTI and Virtual KITTI 2, two datasets with extremely similar scenarios, our approach is capable of making accurate distinctions and aligns well with objective realism. Furthermore, unlike our framework, these prior methods were not originally designed to account for generalization across unseen domains. For instance, [9] fails to disentangle content from style, making them susceptible to content variations. Similarly, [7] and [8] are highly sensitive to environmental factors such as weather and lighting, restricting their scope primarily to clear urban scenes. In contrast, our work explicitly addresses these limitations by validating the framework on the unseen nuScenes and SHIFT datasets.
@@
Photo- realism Enhance Colortransfer[28] 0.242 0.384
@@
CARLA 2Real Raw(CARLA)[6] 0.193 0.307
@@
IV-B4 Sim-to-real Gap Quantification
The results for sim-to-real methods are presented in Table IV. The smaller the value, the more effective the sim-to-real method. Visualization results for photorealism enhance series are shown in Fig. 4d and for CARLA2Real series are shown in Fig. 4e and Fig. 4f. The enhanced images exhibit lower $\mathit{SEDD}_{1}$ and $\mathit{SEDD}_{2}$ values, indicating that our proposed method effectively quantifies the impact of sim-to-real translation. By providing a continuous quantitative measure, our metric allows researchers to precisely compare the effectiveness of different adaptation strategies (e.g., differentiating the superior performance of EPE over CUT), thereby serving as a definitive benchmark for guiding algorithm selection.
@@ IV-D Hyperparameter Experiment
IV-E Discussion
IV-E1 Hyperparameter and Trade-offs
The sensitivity analysis of the temperature $\tau$ and the balancing factor $\lambda$ presented in IV-D serves to characterize the model’s behavior rather than to locate a singular global optimum. This is because the evaluation indicators, intra-class aggregation and inter-class separation, are inherently adversarial. As illustrated in Table VI, varying $\tau$ and $\lambda$ results in a predictable shift where an improvement in one metric often leads to a slight degradation in the other. However, the overall framework effectively converges to a stable equilibrium. Consequently, the grid search conducted only on the Virtual KITTI benchmark is sufficient. Since the primary goal is to identify a stable operating range where both constraints are reasonably satisfied, rather than to overfit the parameters to maximize a specific metric, fine-grained tuning or exhaustive search is unnecessary. This indicates that the framework is not critically sensitive to precise parameter selection, and the established equilibrium is robust enough to generalize without dataset-specific re-optimization.
IV-E2 Limitations
Despite the strong generalization capabilities demonstrated on unseen datasets such as nuScenes, SHIFT and EPE2Vistas, as well as across distinct weather conditions, we acknowledge that the framework may still encounter limitations when applied to scenarios characterized by extremely high diversity or complexity. For instance, samples with extreme illumination conditions, such as night-time scenes. In such cases, the capacity of the current feature and style extraction modules to fully disentangle style from content may be challenged. Future research could investigate more advanced architectures for feature and style extraction to achieve a more rigorous disentanglement.
@@ V Conclusion
In this work, we propose a novel style-based profiling framework with the SEDD metric. Our approach leverages ResNet-based architectures as the backbone network for extracting effective features from input images. To capture the stylistic characteristics, we employ Gram matrices to generate style embeddings, which are subsequently processed through fully connected neural networks for dimensionality reduction while maintaining essential style information. The proposed metric learning framework utilizes Center Loss and NTXent Loss to optimize the model. Extensive experiments validate the effectiveness of the proposed metric.
This work transforms the synthetic-to-real gap into quantifiable objective indicators, eliminating reliance on subjective experience. It provides a standardized proactive tool for the quality control of synthetic datasets in autonomous driving, facilitating more targeted improvements. Future work could explore leveraging SEDD as an objective function for generative models, creating a closed-loop system that actively guides high-fidelity data synthesis.